write a quadratic equation with the given solutions -9 and -4
Answer by cosin(1)

if x=-9 and x=-4 then we have
x+9=0 and x+4=0
or (x+9)(x+4)=0
or x^2+4x+9x+36=0
x^2+13x+36=0